OpenStudy (ashchu117):

Two tangents are drawn from a point. A, which is 37 cm from the center of the circle. The diameter of the circle is 24 cm. What is the length of each tangent from point A to the point of tangency? A. 13 cm B. 32 cm C. 35 cm D. 44cm

OpenStudy (matlee):

no idea

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Use the Pythagorean theorem: \(x^2 + 12^2 = 37^2\) Solve for x.

OpenStudy (ashchu117):

\[2\sqrt{253}\]

OpenStudy (ashchu117):

@Mathstudent55 is that right?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The legs are x and 12 cm. The hypotenuse is 37 cm. The Pythagorean theorem gives us the equation \(a^2 + b^2 = c^2\) where a and b are the legs, and c is the hypotenuse. First, we substitute a, b, and c with the values of our problem: \(x^2 + (12~cm)^2 = (37~cm)^2\) \(x^2 + 144~cm^2 = 1369~cm^2\) Subtract 144 cm^2 from both sides rto get: \(x^2 = 1225~cm^2\) Take square roots of both sides: \(x = 35~cm\)

OpenStudy (ashchu117):

@mathstudent55 so the answer is C?

OpenStudy (anonymous):

\[\sqrt{37^2-12^2}=35 \]C.

OpenStudy (ashchu117):

@robtobey thanks

OpenStudy (mathstudent55):

@Ashchu117 After I solved the equation, and wrote as my last line above x = 35 cm, that was still not enough for you to figure out the answer is C, so you still need to ask? Since choice C is 35 cm, it looks pretty obvious that C is the answer.